​​ 2sin π/8= √(2 + √2)

​​2sin π/8=√(2 + √2)

Grade:11

1 Answers

ng29
8 years ago
there are several values of trigonometry that are need to be learnt  like 150, 750, 180,360.22.50
if u dont know than u can refer some book which contains all of it
i am writing some of them for u
 
tan15= 2-3(1/2) 
tan 22.50 = 2(1/2) -1
cos 360 = (1+5(1/2)) /4
approve if useful